### v1

# Capturing the Demon in Szilard’s Engine

Abstract
In Szilard’s engine, a demon measures a one-particle gas and applies feedback to extract work from thermal fluctuations, embodying Maxwell’s notion that information reduces thermodynamic entropy - an apparent second-law violation. The Landauer-Bennett Thesis resolves this paradox by requiring the demon to record the measurement, which results in an entropy increase in the demon’s memory. Eventually, the demon’s memory needs to be erased. The erasure costs the same work as extracted previously, hence there is no violation of the second law. Though widely accepted, the fictitious memory invoked in the thesis has drawn multiple criticisms, with debates persisting over the demon’s necessity. We show that the demon is the piston that partitions the space and drives the expansion. The final position of the piston after expansion records the particle’s position pre-expansion: it is an “information-bearing degree of freedom”. In this Piston-Demon Thesis, memory register and feedback (expansion) happen simultaneously. Our exposition identifies the mischievous demon as a physical degree of freedom, and greatly simplifies Szilard’s engine. It also offers educators a tangible illustration of information-thermodynamics.
I Introduction: Szilard’s Engine and Landauer-Bennett Thesis
Szilard’s engine was introduced by Leo Szilard in 1929 [1] to realize the paradoxical Maxwell’s demon in the much simplified model of a single-particle ideal gas. Illustrating vividly the interplay between information and energy, Szilard’s engine has long been deemed as a cornerstone of information-thermodynamics and a staple in physics education and research [4].
Consider a single-particle ideal gas in a box of volume $V$ and in thermal contact with a bath at temperature $T$, as schematized in Fig. 1(a). A demon inserts a partition (i.e., the piston) in the middle of the box, confining the particle to the left or right side with equal probability (50%), arriving at state (b) in Fig. 1. The demon measures the particle location (L or R) and initiates a quasi-static expansion. At the end of the expansion, the piston arrives at the left/right end of the box and the volume of the gas restores to $V$, arriving at state (c) in Fig. 1. The work done by the piston during the expansion stage is
[…]
where $P=T/V^{\prime}$ follows the ideal gas law, and $V^{\prime}$ is the instantaneous volume. Throughout this article, we set the Boltzmann constant to unity and use base-2 logarithms to define entropy. Thus, we express temperature in energy units and entropy in bits, a convention that simplifies discussions of information-thermodynamics. The work (1) comes solely from the internal energy of the bath. By energy conservation, the energy of the heat bath decreases by $T$, and the entropy of the bath decreases by $\Delta S_{\text{bath}}=-1\,{\rm bit}$. This would be a violation of the second law, if the demon’s state remains the same before and after the process.
In the commonly accepted Landauer-Bennettthesis [2, 3, 4], the demon must record the result of his measurement (L or R) in his memory, shown as the small boxes beneath the gas boxes in Fig. 1. It is assumed that the memory starts from its standard state L in state (a) and arrives at L or R in state (b), depending on the measurement result. This memory register results in an entropy increase by 1 bit of the demon, or, more precisely, of its memory. According to the modern analysis of Sagawa and Ueda [5, 6], this memory register is also accompanied by a mutual information between the demon’s memory and the position of the particle. At the end of the expansion the system arives at state (c), where the demon’s memory is de-correlated with the state of the gas. Even though the memory still has entropy $1$ bit, it no longer knows the whereabouts of the particle. The memory encodes the particle’s position in state (b), not in state (c).
Figure 1: Landauer-Bennett interpretation of Szilard’s cycle in three states (a), (b), and (c) with transitions. Each dashed box describes a state of the universe that contains a one-particle gas, a piston (dashed line), and a memory bit (small square below). In (a), memory is standard (L). Insertion (a) $\to$ (b): the demon measures and records the particle’s position. Expansion (b) $\to$ (c): the demon uses this to direct expansion, extracting work $T$. Erasure (c) $\to$ (a): the demon erases its memory by resetting the bit back to state L, dissipating $T$. No net work is extracted, and the second law remains intact. Note: the piston (dashed) remains unreset even after erasure of the bit, an issue that the Landauer-Bennett Thesis overlooks.
The process from (a) $\rightarrow$ (b) $\rightarrow$ (c) is not truly cyclic, because the initial state and the final state of the memory are different. As pointed out by Landauer, to make the process cyclic, one must erase the information stored in the memory, i.e., we need to reset the memory bit to its initial state L. According to Landauer and Bennett, this erasure is logically irreversible and requires work of $T$ to be dissipated into the bath. The full process (a) $\rightarrow$ (b) $\rightarrow$ (c) $\rightarrow$ (a), which consists of insertion, expansion, and erasure, is genuinely cyclic and respects the Kelvin-Planck statement of the second law: “It is impossible to devise a cyclically operating device, the sole effect of which is to absorb heat from a single reservoir and deliver an equivalent amount of work”. Indeed, the work extracted from the heat bath during expansion is returned to the heat bath during the erasure. The total entropy of the universe does not change, since the bath, the gas, and the demon all returned to their initial state. The Szilard cycle is thermodynamically reversible.
II Criticisms of Landauer-Bennett Thesis
While the Landauer-Bennettthesis is widely accepted, the fictitious nature of the demon and its memory have been a weak spot that incurred multiple attacks. Feyerabend and Popper proposed mechanical setups that can be used to extract work without measurement [12]. Similarly, Jauch and Baron emphasized that the potential for work exists pre-measurement, rooted in physical constraints rather than subjective knowledge [13]. Alicki [16] reinforced this idea by showing, in his Model B, that the process of work extraction can be started autonomously, i.e., without any demon. Earman and Norton [10, 11] argued that it is possible to design a cycle such that memory erasure is no longer needed. Norton [9] argued that work extraction hinges on a massive piston instead of recording/erasing a bit. According to these works, the expansion and work extraction can be accomplished without the intervention of any demon. Yet these works did not explain how the second law of thermodynamics could be saved in the absence of a demon. In the setting of general information-processing thermodynamics, Bub [14] and Bennett [15] argued that certain variables must be “information-bearing”, and therefore may act like the demon’s memory. Yet, as far as we can see, no attempt has been tried to identify such a concrete information-bearing variable in Szilard’s engine - the demon remains floating in the air. At the very least, Landauer-Bennettthesis needs to be completed by a concrete description of how the demon does his job.
Smoluchowski [7] constructed a fully mechanical trap-door model that may mimic the action of a demon but with no dubious measurement/feedback step. He showed that the system eventually achieves equilibrium with the environment, where there is no possibility of work extraction. The same result was also demonstrated by Feynman using his famous ratchet model [8]. The models of Smoluchowski and Feynman are stochastic and autonomous, the latter means that the systems evolve according to a given Hamiltonian and do not require intervention of any demon. These systems eventually reach equilibrium, and detailed balance (or microscopic reversibility, as called by Onsager) of thermal equilibrium prohibits any work extraction. Yet according to the modern theory of stochastic thermodynamics [17], work extraction in autonomous systems is possible, as long as the systems start from a low entropy initial non-equilibrium state. One vivid illustration of this possibility is the toy model of autonomous demon due to Mandal and Jarzynski [18]. In this model, a binary memory tape is explicitly introduced. If the initial state of the memory is in equilibrium with the heat bath, however, the work would be identically zero. Hence the works of Smoluchowski and Feynman do not exclude the possibility of a demon doing his job. But the question remains: how does the demon do his job?
III The Piston-Demon Thesis
Figure 2: Piston-Demon interpretation of Szilard’s cycle in three states (a), (b), and (c) with transitions. Each dashed box describes a state of the universe that contains a one-particle gas and a piston (filled rectangle) - no demon is needed with this autonomous setup. Insertion (a) $\to$ (b), no recording occurs. Expansion (b) $\to$ (c): extraction work $T$, and recording of the past position of the particle before expansion. Reset (c) $\to$ (a): resetting the piston to the initial position dissipates $T$. No net work is extracted, and entropy remains balanced autonomously. The second law remains intact.
We agree with Feyerabend and Popper, Jauch and Baron, Earman and Norton, and Alicki, and many others, that there is no need to invoke a demon to start the expansion. Indeed, from the perspective of thermodynamics, the potential for work extraction lies in the low entropy nature of the gas state right after the insertion. The tendency of gas expansion is precisely the tendency for an isolated system to spontaneously converge to the entropy-maximizing equilibrium state (the zeroth law of thermodynamics).
Our key observation is in Landauer-Bennett Thesis, the piston is not reset back to its initial state, even though the fictitious memory has been. Hence Szilard’s cycle in Landauer-Bennett Thesis is not truly cyclic. A careful inspection shows that the piston, which drives the expansion, also acts as a memory register, whose final position after expansion records the direction of expansion and therefore also the initial position of the particle. If the particle is on the left, the piston expands rightward to the right wall; if on the right, leftward to the left wall, see Fig. 2 (c). The piston is the “information-bearing degree of freedom” as discussed by Bub [14] and Bennett [15]. Our observation fully reconciles the idea of memory register with the idea of many Landauer-Bennett’s objectors that no measurement/memory register is needed before work extraction. According to our Piston-Demon Thesis, the process of expansion is also the process of measurement/feedback: they are the one and same thing. This should be strongly contrasted with the Landauer-Bennett Thesis that measurement and feedback happen sequentially. Our Piston-Demon Thesis eliminates the need for any fictitious demon - the piston is the demon.
After the expansion, we must reset the piston to its initial position to make the process truly cyclic. This reset is logically irreversible and requires work of $T$ to be dissipated into the bath. The full process (a) $\rightarrow$ (b) $\rightarrow$ (c) $\rightarrow$ (a), as shown in Fig. 2 is cyclic, and there is no change in the total entropy of the universe.
The Piston-Demon Thesis refines and simplifies the Landauer-Bennett Thesis. It combines the roles of the demon and the piston into a single, physical degree of freedom; it condenses memory register and feedback into a single process. It greatly simplifies the logical structure of information-thermodynamics interplay, and makes the Landauer-Bennett Thesis more consistent.
III.1 Entropy balance sheets of two theses
We supply a more detailed comparison between Landauer-Bennett Thesis and Piston-Demon Thesis by studying their entropy balances. In Piston-Demon Thesis, we use $S(G),S(P)$ to denote the entropies of the gas and the piston, and use $I(G;P)$ to denote their mutual information. Their joint entropy is then:
$S(GP)=S(G)-I(G;P)+S(P).$ (2)
The initial entropy of the gas in (a) is $S(G)=\log V/\lambda_{T}^{3}$, where $\lambda_{T}$ is the thermal de Broglie wavelength. Note that $S(G)$ and $S(P)$ are unconditional entropies. If we knew that the particle is on the left handside of the piston after insertion, the conditional entropy of the gas given the information would be $\log(V/2\lambda_{T}^{3})=\log(V/\lambda_{T}^{3})-1$. But that conditional entropy is irrelevant in Piston-Demon Thesis - the piston does not measure the particle right after the insertion. Instead, we infer that information after the expansion is done.
We use $S(B)$ to denote the thermodynamic entropy of the bath. We assume that there is no correlation between the bath and the gas/piston. Hence the total entropy of the universe in Piston-Demon Thesis is
$S_{\rm tot}=S(G)-I(G;P)+S(P)+S(B).$ (3)
We use $S_{0}$ to denote the value of bath entropy in the initial state (a) of the cycle. Breaking down of the total entropy into various components in each of states (a), (b), (c) are listed in Table 1. Note that the total entropy remains the same in all states, consistent with the reversible nature of Szilard’s cycle.
State $S(G)$ $I(G;P)$ $S(P)$ $S(B)$ $S_{\text{tot}}$
(a) $\log_{2}(V/\lambda_{T}^{3})$ 0 0 $S_{0}$ $\log_{2}(V/\lambda_{T}^{3})+S_{0}$
(b) $\log_{2}(V/\lambda_{T}^{3})$ 0 0 $S_{0}$ $\log_{2}(V/\lambda_{T}^{3})+S_{0}$
(c) $\log_{2}(V/\lambda_{T}^{3})$ 0 1 $S_{0}-1$ $\log_{2}(V/\lambda_{T}^{3})+S_{0}$
(a) $\log_{2}(V/\lambda_{T}^{3})$ 0 0 $S_{0}$ $\log_{2}(V/\lambda_{T}^{3})+S_{0}$
Table 1: Entropy balance sheet in the Piston-Demon Thesis for Szilard’s cycle, which is illustrated in Fig. 2. $S(G)$, $S(P)$, and $S(B)$ are entropies of the gas, piston, and bath; $I(G;P)$ is mutual information between current gas and piston states (0 bits at (c), though piston correlates with past gas state); $S_{\text{tot}}$ is total entropy. Note that the total entropy remains invariant, showing the reversibility of Szilard’s cycle.
In Landauer-Bennett Thesis, we use $S(D)$ to denote the entropy of the demon, and use $I(G;D)$ to denote its mutual information with the gas. The joint entropy of the gas and the demon is then:
$S(GD)=S(G)-I(G;D)+S(D).$ (4)
The total entropy of the universe in Piston-Demon Thesis is
$S_{\rm tot}=S(G)-I(G;D)+S(D)+S(B).$ (5)
Breaking down of the total entropy into various components in each of states (a), (b), (c) are listed in Table 2. Note that the total entropy remains the same in all states, consistent with the reversible nature of Szilard’s cycle.
State $S(G)$ $I(G;D)$ $S(D)$ $S(B)$ $S_{\text{tot}}$
(a) $\log_{2}(V/\lambda_{T}^{3})$ 0 0 $S_{0}$ $\log_{2}(V/\lambda_{T}^{3})+S_{0}$
(b) $\log_{2}(V/\lambda_{T}^{3})$ 1 1 $S_{0}$ $\log_{2}(V/\lambda_{T}^{3})+S_{0}$
(c) $\log_{2}(V/\lambda_{T}^{3})$ 0 1 $S_{0}-1$ $\log_{2}(V/\lambda_{T}^{3})+S_{0}$
(a) $\log_{2}(V/\lambda_{T}^{3})$ 0 0 $S_{0}$ $\log_{2}(V/\lambda_{T}^{3})+S_{0}$
Table 2: Entropy balance sheet in the Landauer-Bennett Thesis for Szilard’s cycle, which is illustrated in Fig. 1. $S(G)$, $S(D)$, and $S(B)$ are entropies of the gas, demon, and bath; $I(G;D)$ is mutual information between current gas and demon (1 bit at (b), before expansion); $S_{\text{tot}}$ is total entropy. Note that according to Landauer-Bennett, a mutual information between G and D is established in state (b), before expansion. Note that the total entropy remains invariant, showing the reversibility of Szilard’s cycle.
Now comparing Table 2 with Table 1, we see that besides the nominal difference between D and P, the only difference is in state (b) (after insertion but before expansion): whereas in Landauer-Bennett Thesis, there is a mutual information 1 bit between the gas and the demon, there is no mutual information between the gas and the piston in state (b). This difference reflects the essential difference between these two theses. In Landauer-Bennett Thesis, measurement is done right after the insertion, and expansion (feedback) follows sequentialy. Therefore, in state (b), there is already mutual correlation between the gas and the demon. By contrast, in Piston-Demon Thesis, measurement and feedback happen simultaneously. The demon (actually us) infer the whereabouts of the particle at the time of insertion now instantaneously, but in a later time, when the expansion is done. We believe that this is a universal feature of all information-thermodynamic models when the demon is concretized as physical, information-bearing degree of freedom: measurement and feedback always happen simultaneously.
IV Conclusion
Our Piston-Demon Thesis of the Szilard’s engine eliminates the need for a fictitious demon and makes Landauer’s principle more tangible. It provides a self-contained, intuitive framework for teaching and research of information-thermodynamics.
The author thanks Grok 3, an AI assistant developed by xAI, for both enlightening scientific discussions and useful editorial assistance. The authors have no conflicts to disclose.